calculate the size of the multiplier
1
Expert's answer
2020-09-28T07:53:49-0400

In economics a multiplier is a factor of proportionality that measures the rate at which endogenous variables change with changes in exogenous variables.

How to calculate multiplier.

multiplier is calculates by dividing one over The marginal propensity to save plus the marginal rate of tax on income plus the marginal propensity to import goods and services.

that is

Multiplier("\alpha" ) = 1 / {(sum of the propensity to save("\iota" ) + tax("\Gamma" ) + import("\chi" ) }

"\alpha" =1 above

"\alpha=1/\iota+\Gamma+\chi"